# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

### v4

@@ 7 Conclusion
In this work, we introduced PnP-CoSMo, a modular approach to multi-contrast reconstruction which combines content/style modeling with iterative reconstruction, offering the reconstruction quality of state-of-the-art end-to-end methods with improved cross-contrast generalizability despite significantly lenient training data requirements. At its core is the content consistency operation, which provides regularization at the semantic level of the image content. We defined two quantities that determine the efficacy of this operation, namely optimal content capacity and content discrepancy, and provided several ways of maximizing this efficacy. On real-world clinical data, PnP-CoSMo provided up to 32.6% more acceleration over PnP-CNN for given SSIM, enabling sharper reconstructions at high accelerations. In our preliminary radiological task of visual quality assessment and brain metastasis diagnosis at realistic accelerations, PnP-CoSMo produced images of at least fairly diagnostic overall quality at =5, enabling 60% more acceleration over clinical reconstructions. To progress from our current proof-of-concept towards a practical implementation, future work will focus mainly on improving and extending the content/style model, reducing reconstruction latency, tackling the problem of model hallucinations, incorporating online registration, and conducting a large-scale radiological study.

### v5

@@ 7 Conclusion
In this work, we introduced PnP-CoSMo, a modular approach to multi-contrast reconstruction which combines content/style modeling with iterative reconstruction, offering the reconstruction quality of state-of-the-art end-to-end methods with improved cross-contrast generalizability despite not requiring any k-space training data. At its core is the content consistency operation, which provides regularization at the semantic level of the image content. We defined two quantities that determine the efficacy of this operation, namely optimal content capacity and content discrepancy, and provided several ways of maximizing this efficacy. On real-world clinical data, PnP-CoSMo provided up to 32.6% more acceleration over PnP-CNN for given SSIM, enabling sharper reconstructions at high accelerations. In our preliminary radiological task of visual quality assessment and brain metastasis diagnosis at realistic accelerations, PnP-CoSMo produced images of at least fairly diagnostic overall quality at =5, enabling 60% more acceleration over clinical reconstructions. To progress from our current proof-of-concept towards a practical implementation, future work will focus mainly on improving and extending the content/style model, reducing reconstruction latency, tackling the problem of model hallucinations, incorporating online registration, and conducting a large-scale radiological study.